OpenStudy (anonymous):

What is the gravitational field strength at a height h above the surface of the Earth? R is the radius of the earth

OpenStudy (anonymous):

What is the orbital radius and speed of a synchronous satellite which orbits the earth once every 24h? Take G=66710−11Nm2kg2, Mass of the earth is 5981024kg

OpenStudy (anonymous):

The gravitational field strength of a fixed mass \(m\), \(h\) meters above the Earth's surface is: \[F(h) = G\frac{mM}{(R+h)^2}\] where \(M\) is the mass of the Earth (in \(kg\)) and \(R\) is its average radius (in meters). When dealing with a very large body of mass \(M\) (as the Earth) and another object \(m \) is orbiting it, it means that the gravitational force between them is providing the centripetal force that makes the object \(m\) follow a circular path around \(M\): \(\large{F_{centripetal} = F_{gravitational}}\) \(\large{ma_{c} = G\frac{mM}{r^2}}\) \(\large{mr\omega^2 = G\frac{mM}{r^2}}\) \(\large{\omega^2 = G\frac{M}{r^3}}\) Where \(\omega\) is the angular velocity. The angular velocity is the time it takes the satellite to complete one revolution: \(\frac{2\pi}{T}\), \(T\) is the time it takes to do so. Remember it must be expressed in \(rad/s\). With this information you can solve for the orbital radius \(r\) and then find its tangential velocity by using the relationship between \(v\) and \(\omega\).